# It Is Giving Major Satisfaction: Why Fairness Matters for Software Practitioners

### v3

@@ 3.1. Theoretical Constructs
Fairness Perceptions. We adopt Colquitt’s multidimensional framework of organizational fairness (Colquitt, 2001), which defines fairness across four dimensions: distributive, procedural, interpersonal, and informational fairness.
Job Satisfaction. To assess job satisfaction in the SE context, we draw on the framework by Storey et al. (Storey et al., 2021), which outlines 44 job satisfaction factors relevant to software developers.
Fairness Experiences in SE. We also explore fairness perceptions across specific SE scenarios. Drawing on prior work by Sesari et al. (Sesari et al., 2024), we included items that capture the extent of fairness challenges in SE, such as unfair allocation of work, unfair income distribution, and unfair recruitment.
@@ 3.2. Framing the Research Questions
Based on prior organizational behavior research, we theorize that fairness perceptions influence job satisfaction. Colquitt’s meta-analytic review (Colquitt et al., 2001) found that perceptions of distributive, procedural, interpersonal, and informational fairness each significantly predict job satisfaction across industries (Judge and Colquitt, 2004; Konovsky and Cropanzano, 1991; Moorman, 1991). Drawing on these findings and emerging SE research on fairness (German et al., 2018; Sesari et al., 2024), we posit that developers who perceive higher fairness across these dimensions will report higher job satisfaction. Thus, we formulate the following research questions:
@@
While RQ1 explores the general relationship between fairness perceptions and job satisfaction, existing literature also highlights demographic disparities in SE, a field predominantly young, White, and male. Studies have shown that minority groups often face unique challenges. For example, Trinkenreich et al. (Trinkenreich et al., 2022a) identified that women in the software industry encounter issues such as work-life balance, the glass ceiling, and sexism. Canedo et al. (Canedo et al., 2019) found that women are underrepresented in software development and open source projects, often due to male-dominated workplace conditions. Nadri et al. (Nadri et al., 2021) reported that less than 10% of contributions to open source projects come from visibly non-White developers. Similarly, Baltes et al. (Baltes et al., 2020) discussed age-related biases, noting that older developers are often perceived as less adaptable, which can hinder career progression.
Given these disparities, we theorize that the majority’s perspective may not fully capture the experiences of minority groups. By examining perceptions across diverse demographics, we can ensure that our findings reflect a broader range of experiences.
@@
Building on RQ2 and RQ3, we also explore how the relationship between fairness perceptions and job satisfaction in SE contexts varies across demographic groups. For instance, if respectful treatment during evaluations affects job satisfaction, which demographic groups are most impacted? Is this effect stronger for female developers, suggesting gender plays a critical role?
@@ 4. Methodology
Finally, we detail the steps employed to answer the research questions posed in this study. The scripts for all steps are provided in the supplementary material (fig, 2024).
@@ 4.2. Pre-processing and Assessing Survey Data Quality
To ensure the validity and reliability of our survey constructs, we conducted both Exploratory Factor Analysis (EFA) and Confirmatory Factor Analysis (CFA), followed by an assessment of internal consistency using the Omega reliability coefficient (McDonald, 1970).
@@ 4.4. Grouping Job Satisfaction Factors
After establishing the relationships between factors through correlation analysis, we computed the dissimilarity matrix using the Manhattan distance, which is appropriate for ordinal data. The Manhattan distance measures the absolute differences between the pairwise correlation coefficients, summing these differences across all pairs (Weiß, 2019). This method helps to quantify how dissimilar each pair of job satisfaction factors is based on their correlation values (Weiß, 2019). For instance, if the Manhattan distance between satisfaction with benefits and satisfaction with salary is low, it suggests these factors are closely related in how respondents perceive them.
We then performed hierarchical clustering using the average linkage method, implemented in R with the hclust() function (Langfelder and Horvath, 2012; Guénard and Legendre, 2022). The average linkage method calculates the average distance between all pairs of factors in two clusters, merging the clusters with the smallest average distance first. The result of the hierarchical clustering is visualized in a dendrogram (see Figure 1), a tree-like diagram that illustrates how clusters are merged at different distance thresholds. For instance, if satisfaction with benefits and satisfaction with salary are closely related, they will be merged early in the dendrogram.
@@ 5.3. RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
Takeaway 1: Fairness across all four dimensions; distributive, procedural, interpersonal, and informational, significantly enhances software practitioners’ overall job satisfaction and satisfaction with job security.
@@ 8. Conclusion
Organizations should focus on applying all four fairness dimensions, particularly interpersonal fairness, within SE contexts like authorship, while tailoring practices to meet the needs of diverse demographic groups. Specific strategies have been outlined to address fairness concerns for software practitioners collectively (see Section 6.16.3) and based on their backgrounds (see Section 6.46.7). Future research could explore remedies for fairness problems that have already occurred, examine how fairness manifests in SE contexts, evaluate the effectiveness of suggested strategies, and bridge the gap between collective and demographic-specific fairness concerns.

### v4

@@ 1. Introduction
In this study, our primary objective is to examine how software practitioners’ perceptions of fairness influence their job satisfaction. The main contributions of this paper are as follows: (1) we empirically demonstrate that fairness perceptions are significantly associated with job satisfaction among software practitioners; (2) we show that this relationship differs across demographic groups; and (3) we identify which fairness dimensions matter the most in specific SE contexts and for which demographic groups.
@@ 3.1. Theoretical Constructs
Fairness Perceptions. We draw on Colquitt’s framework of organizational justice (Colquitt, 2001), which defines fairness along four dimensions: distributive fairness (perceived fairness of outcomes, such as rewards or recognition), procedural fairness (fairness of the processes used to determine those outcomes), interpersonal fairness (respectful and dignified treatment by others), and informational fairness (honesty and clarity in communications). These dimensions have been shown to significantly predict job satisfaction across various organizational settings (Judge and Colquitt, 2004; Konovsky and Cropanzano, 1991; Moorman, 1991). In our study, we conceptualize fairness perception as developers’ subjective evaluation of how fairly they are treated across these four dimensions. These perceptions are treated as predictors of job satisfaction.
Job Satisfaction. Job satisfaction reflects the extent to which individuals feel fulfilled, valued, and content with their work. In SE, this is influenced by a combination of social and technical factors. Storey et al. (Storey et al., 2021) identified 44 job satisfaction factors specific to software practitioners, including appreciation, autonomy, work-life balance, and rewards. In this study, we treat job satisfaction as the main outcome variable. By doing so, we aim to examine how perceived fairness contributes to job satisfaction.
Fairness Experiences in SE Contexts. While fairness perceptions are often assessed at a general level, we also examine how they are experienced in specific SE scenarios. Building on Sesari et al. (Sesari et al., 2024), who identified key fairness contexts in SE, such as work allocation, recruitment, income, and evaluations, we include items that capture developers’ fairness experiences across these contexts. This allows us to explore not just whether fairness matters, but also when it matters the most.
@@ 3.2. Framing the Research Questions
This question is grounded in prior organizational behavior research, which suggests that fairness perceptions influence job satisfaction. Colquitt’s meta-analytic review (Colquitt et al., 2001) found that perceptions of distributive, procedural, interpersonal, and informational fairness each significantly predict job satisfaction across industries (Judge and Colquitt, 2004; Konovsky and Cropanzano, 1991; Moorman, 1991). Drawing on these findings and emerging SE research on fairness (German et al., 2018; Sesari et al., 2024), we posit that developers who perceive higher fairness across these dimensions will report higher job satisfaction.
@@
While RQ1 explores the general relationship, existing literature highlights demographic disparities in SE, a field predominantly young, White, and male. Studies have shown that minority groups often face unique challenges. For example, Trinkenreich et al. (Trinkenreich et al., 2022a) identified that women in the software industry encounter issues such as work-life balance, the glass ceiling, and sexism. Canedo et al. (Canedo et al., 2019) found that women are underrepresented in software development and open source projects, often due to male-dominated workplace conditions. Nadri et al. (Nadri et al., 2021) reported that less than 10% of contributions to open source projects come from visibly non-White developers. Similarly, Baltes et al. (Baltes et al., 2020) discussed age-related biases, noting that older developers are often perceived as less adaptable, which can hinder career progression. Given these disparities, we theorize that the majority’s perspective may not fully capture the experiences of minority groups. By examining perceptions across diverse demographics, we can ensure that our findings reflect a broader range of experiences.
@@
Building on our findings from RQ2, where we observe whether the relationship between fairness perceptions and job satisfaction varies across demographic groups, it becomes essential to explore this dynamic in specific SE scenarios. Given the insights from RQ3, we also examine how the manifestations of fairness issues differ across demographic groups within SE contexts. For instance, if we find that being treated politely, with respect, and dignity during evaluations significantly affects job satisfaction, we should investigate which demographic groups this applies to the most. Is it more impactful for female developers, suggesting that gender plays a key role in this relationship?
@@ 4. Methodology
Finally, we detail the steps employed to answer the research questions posed in this study. The scripts for all steps are provided in the supplementary material (fig, 2024) and Fig. 1 illustrates the outline of our methodology.
Figure 1. Overview of Methodology
@@ 4.2. Pre-processing and Assessing Survey Data Quality
To ensure the validity and reliability of our survey constructs, we conducted both Exploratory Factor Analysis (EFA) and Confirmatory Factor Analysis (CFA), followed by an assessment of internal consistency using the Omega reliability coefficient (McDonald, 1970) (see Table 3).
Table 3. Model Fit, Factor Loadings, and Reliability by Construct
Analysis Metric Fairness Perception Job Satisfaction Fairness Context in SE
EFA RMSEA 0.06 0.08 0.08
TLI 0.94 0.87 0.90
RMSR 0.04 0.05 0.07
CFA CFI 0.92 0.86 0.92
TLI 0.91 0.83 0.90
RMSEA 0.08 0.10 0.09
SRMR 0.08 0.08 0.06
Factor Loadings (range) 0.52–0.97 0.49–0.87 0.41–0.78
Omega Reliability 0.94 0.93 0.89
@@ 4.4. Grouping Job Satisfaction Factors
After establishing the relationships between factors through correlation analysis, we computed the dissimilarity matrix using the Manhattan distance, which is appropriate for ordinal data. Compared to the Euclidean distance, the Manhattan distance provides a more robust measure when dealing with ordinal or non-normally distributed variables, as it does not assume interval-scaled data or linear relationships (Řezanková and Novák, 2019). The Manhattan distance measures the absolute differences between the pairwise correlation coefficients, summing these differences across all pairs (Weiß, 2019). This method helps to quantify how dissimilar each pair of job satisfaction factors is based on their correlation values (Weiß, 2019). For instance, a low value for the Manhattan distance between satisfaction with benefits and satisfaction with salary suggests that these factors are closely related in how respondents perceive them.
We then performed hierarchical clustering using the average linkage method, implemented in R with the hclust() function (Langfelder and Horvath, 2012; Guénard and Legendre, 2022). The average linkage method calculates the average distance between all pairs of factors in two clusters, merging the clusters with the smallest average distance first. This approach ensures that clusters are formed based on the overall similarity of their members, rather than just the closest or farthest points, which helps maintain balanced clusters. The result of the hierarchical clustering is visualized in a dendrogram (see Figure 2), a tree-like diagram that illustrates how clusters are merged at different distance thresholds. For instance, if satisfaction with benefits and satisfaction with salary are closely related, they will be merged early in the dendrogram.
@@ 5.3. RQ1: Does fairness perception shape job satisfaction among software practitioners? If so, how?
Takeaway 1: Fairness across all four dimensions (namely distributive, procedural, interpersonal, and informational) significantly enhances software practitioners’ overall job satisfaction and satisfaction with job security.
@@ 8. Conclusion
Organizations should focus on applying all four fairness dimensions, particularly interpersonal fairness, within SE contexts like authorship, while tailoring practices to meet the needs of diverse demographic groups. Specific strategies have been outlined to address fairness concerns for software practitioners collectively (see Section 6.16.3) and based on their backgrounds (see Section 6.46.7). While these findings offer valuable insights, they should be interpreted with caution due to the study’s reliance on self-reported perceptions and its sampling frame as discussed in Section 7. Future research could explore remedies for fairness problems that have already occurred, examine how fairness manifests in SE contexts, evaluate the effectiveness of suggested strategies, and bridge the gap between collective and demographic-specific fairness concerns.